# Nonlocal Retinex-Based Variational Model and its Deep Unfolding Twin for Low-Light Image Enhancement

### v1

@@ 3.1 Definitions and notations
$\left(\sum_{i=1}^{n_{i}}\left(\sum_{j=1}^{n_{j}}\left(\sum_{k=1}^{n_{k}}\left(\sum_{t=1}^{n_{t}}\left|x_{i,j,k,t}\right|^{s_{t}}\right)^{s_{k}/s_{t}}\right)^{s_{j}/s_{k}}\right)^{s_{i}/s_{j}}\right)^{1/s_{i}}.$
@@ 3.4 Nonlocal weights
$d\big{(}\tilde{I}_{:,i},\tilde{I}_{:,j}\big{)}=\sum_{z\in B(0,\kappa)\cap\mathbb{Z}}|\tilde{I}_{:,i+z}-\tilde{I}_{:,j+z}|^{2}$
@@
if $j\in B(i,\hat{\nu})\cap\mathbb{Z}$, and zero otherwise. Now, $\hat{\nu}\in\mathbb{Z}^{+}$ determines the size of the search window, $\hat{h}_{\text{sim}}>0$ measures how fast the weights decay with increasing dissimilarity between patches, $\hat{\Gamma}_{i,t}$ is the normalizing factor, and
$d\big{(}(\nabla\hat{I})_{:,i,t}(\nabla\hat{I})_{:,j,t}\big{)}=\sum_{z\in B(0,\hat{\kappa})\cap\mathbb{Z}}|(\nabla\hat{I})_{:,i+z,t}-(\nabla\hat{I})_{:,j+z,t}|^{2}$
@@ 3.6 Automatic gamma correction
$\gamma^{n+1}=\gamma^{n}-\frac{F(\gamma^{n})}{F^{\prime}(\gamma^{n})}=\gamma^{n}-\frac{{\frac{1}{M}}\sum_{i=0}^{M}L_{i}^{\gamma}-0.5}{{\frac{1}{M}}\sum_{i=0}^{M}\left(L_{i}^{\gamma}\ln\left(L_{i}\right)\right)}.$
@@ 4.2 Network architectures
where $Up$ is a pixel-shuffle operation that recovers the original image dimension, and $\widecheck{J}_{P}$ and ˇ$\widecheck{\mathfrak{R}}_{P}^{n}$ are obtained by applying IN and the Linear operator to the projection of $J$ and $\mathfrak{R}^{n}$. The corresponding parameters are shared neither among them nor between the different heads. Finally, following the approach in [24], we repeat the proposed the cross-attention module in the proximity of the nonlocal regularizer by previously shifting $S/2$ the resulting output and $J_{P}$ to avoid artifacts in the border of the image. We have indicated this process in Figure 1 by Patch Cross-Attention (P-CA) and Shifted-Patch Cross-Attention (SP-CA).
@@ 5 Experimental Results
$\text{Loss}\left(I_{out},I_{gt}\right)$ $=\text{MSE}\left(I_{out},I_{gt}\right)+\alpha_{1}\text{Loss}_{c}\left(I_{out},I_{gt}\right)$
$+\alpha_{2}\text{LPIPS}\left(I_{out},I_{gt}\right),$
@@
$\text{Loss}_{c}\left(I_{out},I_{gt}\right)=\frac{\sum_{i=1}^{M}\sum_{k=1}^{C}\left(I_{out}\right)_{k,i}\left(I_{gt}\right)_{k,i}}{MC}.$
@@ 5.3 Experiments on LIME dataset
Output without $N$
Output with $N$
@@ 7 Conclusions
Acknowledgements This work is part of the MoMaLIP project PID2021-125711OB-I00 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGeneration EU/PRTR. Daniel Torres is also supported by the Conselleria d’Educació i Universitats del GOIB through grant FPU2024-002-C.

### v2

@@ 3.1 Definitions and notations
$\left(\hskip-1.9919pt\sum_{i=1}^{n_{i}}\hskip-1.9919pt\left(\hskip-1.9919pt\sum_{j=1}^{n_{j}}\hskip-1.9919pt\left(\hskip-1.42271pt\sum_{k=1}^{n_{k}}\hskip-2.27626pt\left(\sum_{t=1}^{n_{t}}\left|x_{i,j,k,t}\right|^{s_{t}}\hskip-0.85355pt\right)^{s_{k}/s_{t}}\hskip-0.85355pt\right)^{s_{j}/s_{k}}\hskip-0.85355pt\right)^{s_{i}/s_{j}}\hskip-0.85355pt\right)^{1/s_{i}}\hskip-4.26773pt.$
@@ 3.4 Nonlocal weights
$d\big{(}\tilde{I}_{:,i},\tilde{I}_{:,j}\big{)}=\hskip-2.84544pt\sum_{z\in B(0,\kappa)\cap\mathbb{Z}}\hskip-2.84544pt|\tilde{I}_{:,i+z}-\tilde{I}_{:,j+z}|^{2}$
@@
[…]
$d\big{(}(\nabla\hat{I})_{:,i,t},(\nabla\hat{I})_{:,j,t}\big{)}=\hskip-14.22636pt\sum_{z\in B(0,\hat{\kappa})\cap\mathbb{Z}}\hskip-4.26773pt|(\nabla\hat{I})_{:,i+z,t}-(\nabla\hat{I})_{:,j+z,t}|^{2}$
@@ 3.6 Automatic gamma correction
$\gamma^{n+1}=\gamma^{n}-\frac{F(\gamma^{n})}{F^{\prime}(\gamma^{n})}=\gamma^{n}-\frac{{\frac{1}{M}}\sum_{i=1}^{M}L_{i}^{\gamma}-0.5}{{\frac{1}{M}}\sum_{i=1}^{M}\left(L_{i}^{\gamma}\ln\left(L_{i}\right)\right)}.$
@@ 4.2 Network architectures
where $Up$ is a pixel-shuffle operation that recovers the original image dimension, and $\widecheck{J}_{P}$ and $\widecheck{\mathfrak{R}}_{P}^{n}$ are obtained by applying IN and the Linear operator to the projection of $J$ and $\mathfrak{R}^{n}$. The corresponding parameters are shared neither among them nor between the different heads. Finally, following the approach in [24], we repeat the proposed the cross-attention module in the proximity of the nonlocal regularizer by previously shifting $S/2$ the resulting output and $J_{P}$ to avoid artifacts in the border of the image. We have indicated this process in Figure 1 by Patch Cross-Attention (P-CA) and Shifted-Patch Cross-Attention (SP-CA).
@@ 5 Experimental Results
$\text{Loss}\left(I_{out},I_{gt}\right)$ $\hskip-2.27626pt=\hskip-2.27626pt\text{MSE}\left(I_{out},I_{gt}\right)\hskip-1.42271pt+\hskip-1.42271pt\alpha_{1}\text{Loss}_{c}\hskip-1.42271pt\left(I_{out},I_{gt}\right)$
$\hskip-1.42271pt+\hskip-1.42271pt\alpha_{2}\text{LPIPS}\left(I_{out},I_{gt}\right),$
@@
$\text{Loss}_{c}\hskip-1.42271pt\left(I_{out},I_{gt}\right)=\frac{\sum_{i=1}^{M}\sum_{k=1}^{C}\left(I_{out}\right)_{k,i}\left(I_{gt}\right)_{k,i}}{MC}.$
@@ 5.3 Experiments on LIME dataset
Output without NN
Output with NN
@@ 7 Conclusions
Acknowledgements This work is part of the project TED2021-132644B-I00 funded by MICIU/AEI/10.13039/501100011033 and the European Union NextGeneration EU/PRTR, and also by the MoMaLIP project PID2021-125711OB-I00 funded by MICIU/AEI/10.13039/501100011033 and the European Union NextGeneration EU/PRTR. Daniel Torres is also supported by the Conselleria d’Educació i Universitats del GOIB through grant FPU2024-002-C. The authors gratefully acknowledge the computer resources at Artemisa, funded by the EU ERDF and Comunitat Valenciana and the technical support provided by IFIC (CSIC-UV).